Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-06-04
ISS On-Orbit Status 06/04/08

Flight Day 5 (FD5) of STS-124/1J.

ISS crew work cycle remains unchanged: wake 6:32am EDT; sleep 10:02pm.

Arigato Gozaimasu! Congratulations, JAXA! There is Hope in space! At ~5:09pm EDT, the JPM (Japanese Pressurized Module) of the JEM 'Kibo' laboratory complex was opened and ingressed by Aki Hoshide and Karen Nyberg for the first time, joined later by the rest of the crew who clearly enjoyed the voluminous super laboratory. (Kibo is permanently attached at the Node-2 (Harmony) portside hatch since last night.)

FE-1 Oleg Kononenko performed the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS (Russian Segment) hatchways, including the DC1-to-Soyuz tunnel, and the FGB-to-Node passageway. (This is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).)

After last night's successful gross leak check of the Kibo/Node-2 vestibule by MS1 Karen Nyberg following the berthing of Kibo to the Node-2 port hatch (7:01pm), FE-2 Garrett Reisman this morning terminated the overnight fine leak check, which confirmed successful hermeticity of the linkup.

Afterwards, MS4 Aki Hoshide & Nyberg were the principal crewmembers charged with today's opening up of the Japanese laboratory. Before hatch opening and first ingress in Kibo at 5:09pm, --

Hoshide & Nyberg spent ~6 hrs between them on JPM-to-Node-2 vestibule outfitting, setting up equipment including stringing & connecting power cables for routing electricity to Kibo,
Reisman & Hoshide activated Channel 'B' power to provide initial environmental conditions inside the laboratory for human access; this also allowed transfer of command capability for Kibo from MCC-Houston to JAXA's TKSC (Tsukuba Space Center) in Japan which will activate the A-string systems tomorrow (FD6),
Kononenko replaced the aft NPRV (Negative Pressure Relief Valve) in the Node-2 port hatch with the regular IMV (Intermodule Ventilation) valve, and
Aki completed Kibo vestibule outfitting by installing the nitrogen (N2) supply jumper/hose between Node-2 and JPM.
After ingressing the laboratory wearing PPE (Personal Protective Equipment, i.e., surgical masks plus goggles) for the first few minutes until it was determined that the cabin air was clean of any debris, the crewmembers set up the JPM by -

Completing the purge of the N2 system to remove any contamination from the lines,
Collecting air samples in the JPM using the AK-1M sampler kit before air duct installation,
Disconnecting and removing the FSE (Flight Support Equipment) Accumulator from the TCS LTL (Thermal Control System/Low Temperature Loop),
Reconfiguring the TCA GN2 (Thermal Control Assembly Gaseous Nitrogen) Manual Accumulator by opening its isolation valve,
Transferring & installing two hard dummy panels from the JLP (Japanese Logistics Pressurized Module) and installing them in the JPM in preparation of EPS1 (Electrical Power Systems 1) Rack & WS (Work Station) Rack installation (access to JLP requiring temporary activation of the JLP MKAM/Minimum Keep Alive Monitor fan as usual),
Transferring the JEMRMS (JEM Robotic Manipulator System) Rack from the JLP to the JPM,
Deploying necessary drag-through cables from Node-2 into Kibo,
Removing stowage bags from in front of the EPS2 & DMS2 (Data Management System 2) Racks and stowing them near the JEM AL (Airlock), and
Configuring the Kibo lab with two PFEs (Portable Fire Extinguishers) from Node-1 & FGB stowage, plus one PBA/QDMA (Portable Breathing Apparatus/Quick-Don Mask Assembly), a second PBA/QDMA to be added on FD9 after EVA-3.
FE-1 Kononenko had two hours set aside for working a major IFM (In-flight Maintenance) on the Russian ASU toilet facility in the SM (Service Module), removing the failed MNR-NS gas/liquid separator pump of the facility and replacing it with a new unit delivered on STS-124. As of now, the IFM appears to have been successful: the ASU is working nominally, so far.

Major IFM was also performed by FE-2 Reisman who removed the failed desiccant/sorbent bed #2 (202) of the US CDRA (Carbon Dioxide Removal Assembly) and replaced it with a new unit delivered on STS-124. FE-2-17 Chamitoff assisted in ythe task as per 'dedicated handover'.

CDR Sergey Volkov continued the annual inspection and photo-documentation of window panes in the SM, started earlier (5/28). The observed defects were recorded in image and text files on the RSK1 laptop for subsequent downlink via U.S. OCA assets. (Objective of the inspection, which uses a digital still camera (Nikon D1X w/SB-28DX flash) and voice recorder, today was to assess the pane surfaces on SM windows 2, 3, 5, 8, 9 for any changes (new cavities, scratches, new or expanded old stains or discolorations affecting transparency properties) since the last inspection, performed by Oleg Kotov on 6/26/07. The new assessment will be compared to the earlier observations. Defects are measured with the parallax method which uses eyeball-sighting with a ruler and a right isosceles triangle to determine the defects' size and position with respect to the window's internal surface (parallax being the apparent change in an object's position resulting from changing the observer's position).)

Kononenko serviced the Matryoshka-R (RBO-3-2) radiation payload, which has taken over the ESA/RSC-Energia experiment ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS/ALC) with its Spectrometer (AST) and ALC equipment on DC1 panel 429, performing the periodic test on the AST and its current ALC-948 PCMCIA memory card, using the RSK1 laptop for checking the #948 PCMCIA (Portable Computer Memory Card International Adapter).

CDR Mark Kelly & PLT Ken Ham worked in the Node-1 starboard alcove, installing a check valve-hose for A/L CCAA (Common Cabin Air Assembly) safing.

Volkov set up the TTM-2 and 'Kelvin-Video' batteries for charging for another operational run of the Russian KPT-2 science payload BAR-RM. Charging will be terminated tomorrow (6/5), with data gathering starting Friday (6/6) using the RSE-1 laptop, with downlinking via BSR-TM channel. (Objective of the payload is to experiment with ISS leak detection based on environmental data anomalies (temperature, humidity, and ultrasound emissions) at leak locations. The payload uses a remote infrared thermometer (Kelvin-Video), a thermohygrometer (Iva-6A), a heat-loss anemometer/thermometer (TTM-2), an ultrasound analyzer (AU-01), and a leak detector (UT2-03) to determine physical background signs of loss of ISS pressure integrity which could be indicative of leaks in the working compartments of the station. Measurements are taken in specific zones (13 in SM PkhO and 4 in DC1), both with lights & fans turned on and off. )

Sergey also collected Elektron feed water samples from downstream of the BKO multifiltration unit in a drinking bag (paket dlya napitkov) for return to Earth, to monitor the quality of the water being fed from the KOV EDV container through the BKO to the Elektron-VM oxygen generator.

Oleg removed existing tapes from both VDS VTRs (Video Distribution Subsystem/Video Tape Recorders), labeled them for stowage & return to Houston, then loaded the VTRs with fresh tapes.

At ~9:22am EDT, Volkov & Kononenko linked up with MBI-8 PROFILAKTIKA specialists at TsUP-Moscow via VHF/audio to discuss physical exercise needs for the MBI-8 experiment. (Data yields from the TVIS treadmill exercises have been insufficient for analysis. Since the Russian VB-3 VELO ergometer does not record exercise data, the discussion focused on ways for the crew to measure and downlink adequate VELO exercise data for MBI-8.)

FE-2 Reisman spent about 3 hrs on generic handover activities with Greg Chamitoff (who assumes Garrett's FE-2 position after Shuttle departure), starting with Crew Safety Handovers, using a detailed 'handover book'. (During these periods crewmembers are scheduled together to complete various designated standard tasks for familiarizing the new station resident with procedures, caveats, etc. There are also dedicated handovers for selected critical activities. )

Greg Chamitoff also had time reserved for familiarizing himself with the onboard CMS (Crew Medical Systems) exercise equipment.

Kononenko performed the regular updates of onboard SODF (Station Operations Data File) emergency/warning & procedures books with new material delivered on STS-124. (These include specific Emergency Books deployed in the Lab, Soyuz, SM, A/L (normally in FGB) and Node-2, Warning Books in the Lab, SM and FGB, a Leak Pinpoint Repair Kit in the PMA-1 (Pressurized Mating Adapter 1), plus a POC (Portable Onboard Computer) Book and Printer Cue Cards.)

Volkov had 30 min set aside for reviewing the new DOUG (Dynamic Onboard Ubiquitous Graphics) software setups for the upcoming SSRMS (Space Station Remote Manipulator System) and JEM RMS (Robotic Manipulator System) operations planned for tomorrow (FD6)

With assistance by Reisman, Greg Chamitoff performed the periodic offloading of the Lab CCAA (Common Cabin Air Assembly) dehumidifier's condensate tank, filling a CWC (Contingency Water Container, #1062) with the collected water slated for processing. No samples were required. (Transferred quantity is determined by allowing tank and CWC equalize with each other. Tank quantity stabilizes at a value higher than the neutral point. Waiting time ~30 min.)

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2, TVIS treadmill (CDR, FE-1), RED resistive exercise device (CDR, FE-2) and VELO cycle with bungee cord load trainer (FE-1).

Afterwards, Oleg copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

In the SM, the FE-1 completed the routine maintenance of the SOZh/ECLSS system, including ASU toilet facilities systems/replaceables. Removal & replacement of the failed separator pump is on tomorrow's schedule for Kononenko at ~9:30-11:30am EDT. (Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier.)

CDR Volkov performed the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

FE-2-17 Chamitoff had another 2.5 hrs for himself for general orientation (station familiarization & acclimatization) as is standard daily rule for fresh crewmembers for the first two weeks after starting station residency.

Kononenko was to check in Node-2 to verify proper installation of the CVIU (Common Video Interface Unit) VCP J01 video cap at the S3 video camcorder port that is required for transmitting video from the Shuttle to ISS.

Fossum, Garan, Kelly, Ham, Volkov and Chamitoff conducted a one-hour review of timeline & procedures for tomorrow's EVA-2 by Mike Fossum (EV1) & Ron Garan (EV2).

Later, after completing Airlock Equipment Lock (A/L EL) configuration for the spacewalk, including recharging batteries and checking EMU equipment, Mike & Ron will begin their 'campout' in the 'Quest' A/L, starting mask prebreathe, while configuring EVA tools, at ~8:57pm, then closing hatches and initiating depressurization of the CL (Crewlock) from 14.7 to 10.2 psi. Sleep for them and the ISS crew will commence at 10:02pm, for the Shuttle crew at ~10:32pm. (The overnight Campout (nachalo desaturatsiy = desaturation start) in the A/L CL (Crewlock) for denitrogenation/pre-breathe at 10.2 psi lasts about 8.5 hrs. First, the two spacewalkers will perform PBA (Portable Breathing Apparatus) mask prebreathe for denitrogenation, while readying their tools & equipment, then depress the CL from 14.7 to 10.2 psi for their sleep period, to last until ~6:32am EDT tomorrow. The CL hatch will then be cracked at ~7:12am (i.e., temporarily repressurized) for a hygiene break/with mask prebreathe for Fossum & Garan. Around 8:22am, the hatch will be closed again for EVA preps in 10.2 psi, followed by EMU purge & prebreathe. Afterwards, Mark Kelly will support CL depressurization until egress.)

EVA-2, beginning tomorrow nominally at ~11:32am EDT, will last an estimated 6h 30min, i.e., ending at 6:02pm. PLT Ken Ham will again be IV (Intravehicular) crewmember.
--- EVA-1 main objectives are to:

Install Fwd/Aft JTVEs (JEM Television Equipment, i.e., 2 cameras) for monitoring JEM RMS ops,
Remove 7 thermal covers from each of the JEM RMS' six joints & one end effector,
Prepare the JPM zenith ACBM (Active Common Berthing Mechanism) for JLP relocation on FD7 (remove ACBM cover; release MMOD/Micrometeoroid Orbital Debris shield launch restraints),
Clear up late-discovered MLI (Multi-Layer Insulation) cover anomaly on the JPM zenith CBM (could interfere with the 4 capture latches during JLP relocation/berthing),
Install JPM trunnion and keel pin covers,
Prepare S1 Nitrogen Tank Assembly on S1 truss and spare on ESP-3 (External Stowage Platform 3) for R&R during EVA 3, and
Retrieve ETVCG (External TV Camera Group) camera from CP9 (Camera Port 9) on the left truss.
At ~5:35pm, Greg is scheduled for another regular PMC (Private Medical Conference) via S- & Ku-band audio/video.

Transfers Update: N2 transfer from the Shuttle began today. No O2 transfer is panned for 1J. One filled water CWC (#1098) was transferred to the ISS. As of FD3, middeck resupply transfers were 37% of total, overall cargo transfers 12% of total.

No CEO (Crew Earth Observation) photo targets uplinked for today.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use